OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1

OpenStudy (anonymous):

heyyy

OpenStudy (anonymous):

Multiply by 10 both the sides first..

OpenStudy (anonymous):

A. 69x + 10y = 51 B. -69x + 10y = 51 C. 69x + 10y = -51 D. -69x + 10y = -51

OpenStudy (anonymous):

I said multiply by 10 both the sides.. What will you get??

OpenStudy (anonymous):

ummm ok lollll

OpenStudy (anonymous):

-69

OpenStudy (anonymous):

Ha ha ha ha ha... :)

OpenStudy (anonymous):

B(:

OpenStudy (anonymous):

heyyyy

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

What is the graph of the absolute value inequality? |x + 5| y – 2